# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ 2 $\{0\}$$\{1\}$$\{2\}$
Inductive step: Assume that $v(\alpha)$ was defined for every $\alpha\in\Lambda_{n}$ such that $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v\in PLV^{\prime}(\Lambda_{n})$ up to this stage, for a given $n\geq 0$ (Induction Hypothesis, IH). Let $\alpha=\alpha_{m}$ (for $m\in I_{n+1}$) such that $\alpha\not\in\Lambda$ and $\textsf{c}_{1}(\alpha)=n+1$. Observe that $v(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$. Since $v\in PLV^{\prime}(\Lambda_{n})$ then, for every $\delta\in\Lambda_{n}$ such that $v(\delta)=1$ there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda_{n}}(v,w,\delta)$. Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more that one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$). Recall that $\Lambda_{n}^{m}=\Lambda_{n}\cup\{\alpha_{m}\}$ is closed under subformulas. There are four cases to analyze:
[…]
(2.1) There is $\delta\in\Lambda_{n}$ with $v(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=0$. Define $v(\alpha)=1$. Hence $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\delta}$.
(2.2) There is $\delta\in\Lambda_{n}$ with $v(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=1$. By Lemma 3.20, $v^{\prime}=w^{\delta}_{|\Lambda_{n}^{m}}\in PLV^{\prime}(\Lambda_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda_{n}^{m}}(v^{\prime},w^{\prime\prime},\alpha)$. Define $v(\alpha)=1$. Hence, $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\prime\prime}$.
(2.3) For every $\delta\in\Lambda_{n}$ such that $v(\delta)=1$, $w^{\delta}(\alpha)=2$ for every $w^{\delta}$. Define $v(\alpha)=2$. Clearly, $v\in PLV^{\prime}(\Lambda_{n}^{m})$.
[…]
@@
(3.1) $\alpha=\beta\to\gamma$. Then $v(\beta)=2$ and $v(\gamma)=1$. Define $v(\alpha)=1$. Then $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\gamma}$. Indeed, since $w^{\gamma}(\beta)=2$ and $w^{\gamma}(\gamma)=0$ then $w^{\gamma}(\alpha)=0$.
@@
(3.2) $\alpha=\beta\land\gamma$. Then $v(\beta)=1$ and $v(\gamma)\in\mathcal{D}$ or vice versa. Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ (where $w=w^{\beta}$ or $w=w^{\gamma}$) such that $w(\beta)=0$ or $w(\gamma)=0$ and so $w(\alpha)=0$. Define $v(\alpha)=1$. Then $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\beta}$ or $w^{\alpha}=w^{\gamma}$.
[…]
We have shown in cases (1)-(4) how to extend the domain of $v$ to the additional formula $\alpha_{m}$, for every $\alpha_{m}$ such that $\textsf{c}_{1}(\alpha_{m})=n+1$, showing that $v\in PLV^{\prime}(\Lambda_{n}^{m})$. Consider now the simultaneous extension of $v$ with the values assigned to every $\alpha_{m}$ as prescribed above. We obtain so an extension of $v$ to $\Lambda_{n+1}=\cup_{m}\Lambda_{n}^{m}$. By the very definition of partial’ level valuations it follows that $v\in PLV^{\prime}(\Lambda_{n+1})$, and $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$. This completes the inductive step. ∎
@@
Let $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. Then, $\tilde{v}_{p}\in PV(\Lambda)$. Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$. Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$. Let $\tilde{w}_{p}=w_{|\Lambda}$. By Lemma 3.20, $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$. Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$. By Lemma 3.23, there exists $v^{\prime}\in\mathcal{L}^{\prime}_{S4}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$. Hence, $\mathsf{P}_{\Lambda}(\tilde{v}_{p},v^{\prime},\alpha)$. This shows that
$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$}\big{)}\}.\end{array}$ From this, being $PV(\Lambda)$ finite, it follows that $PLV(\Lambda)=PLV^{\prime}(\Lambda)$. ∎
@@
4 An RNmatrix for IPL
@@
Consider the signature Ω={{¬},{,,}}normal-$\Omega=\{\{\neg\},\{\land,\lor,\rightarrow\}\}$ for intuitionistic propositional logic (IPL). The set of intuitionistic formulas will be denoted by as $For(\Omega)$.
4.1 Gödel-McKinsey-Tarski Box Translation
@@
(Box translation function) Let $f_{\Box}:For(\Omega)\rightarrow For(\Sigma)$ be a function such that:
@@
[…]
It is widely known that this translation is faithful [9].
4.2 Truth tables for intuitionistic logic
We introduce the truth tables for intuitionistic logic based on the presented translation.
4.2.1 Intuitionistic non-deterministic valuations
@@
If $v$ is a valuation in $\mathcal{M}_{S4}$, then, for every $\alpha\in For(\Sigma)$, $v(\Box\alpha)=\Box^{\mathcal{A}_{S4}}(v(\alpha))$. This follows from the fact that the multioperation $\Box^{\mathcal{A}_{S4}}$ is deterministic.

### v4

@@ 2 $\{0\}$ $\{1\}$ $\{2\}$
Let $\alpha=\alpha_{\omega\cdot(n+1)+m}$. If $\alpha\in\Lambda$ then ${\Lambda}_{n}^{m+1}={\Lambda}_{n}^{m}$. Then $v_{n}^{m+1}:=v_{n}^{m}$, and the required hypothesis are obviously satisfied by $v_{n}^{m+1}$. Suppose now that $\alpha\not\in\Lambda$. Observe that $v_{n}^{m+1}(\beta)=v_{n}^{m}(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$. Since $v_{n}^{m}\in PLV^{\prime}({\Lambda}_{n}^{m})$ then, for every $\delta\in{\Lambda}_{n}^{m}$ such that $v_{n}^{m}(\delta)=1$ there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}^{4}_{{\Lambda}_{n}^{m}}(v_{n}^{m},w,\delta)$. Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more than one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$). Recall that ${\Lambda}_{n}^{m}$ is closed under subformulas. There are three cases to analyze:
@@
(1) $\alpha=\#\beta$ such that $^{S4}v_{n}^{m}(\beta)=\{0\}$ (for some $#\in\{\neg,\Box\}$), or $\alpha=\beta\#\gamma$ such that $^{\prime S4}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{0\}$ (for some $#\in\{\land,\vee,\to\}$). Then, define $v_{n}^{m+1}(\alpha)=0$. Observe that $v_{n}^{m+1}\in PLV^{\prime}({\Lambda}_{n}^{m+1})$.
@@
(2) $\alpha=\#\beta$ such that $^{S4}v_{n}^{m}(\beta)\in\mathcal{D}$ (for some $#\in\{\neg,\Box\}$), or $\alpha=\beta\#\gamma$ such that $^{\prime S4}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))\in\mathcal{D}$ (for some $\in\{\vee$$\to\}$). There are three subcases to analyze:
(2.1) There is $\delta\in{\Lambda}_{n}^{m}$ with $v_{n}^{m}(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=0$. Define $v_{n}^{m+1}(\alpha)=1$. Hence $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$ such that $w^{\alpha}=w^{\delta}$.
(2.2) There is $\delta\in\Lambda_{n}^{m}$ with $v_{n}^{m}(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=1$. By Lemma 6.13, $v^{\prime}=w^{\delta}_{|\Lambda_{n}^{m}}\in PLV^{\prime}(\Lambda_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}^{4}_{\Lambda_{n}^{m}}(v^{\prime},w^{\prime\prime},\alpha)$. Define $v_{n}^{m+1}(\alpha)=1$. Hence, $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$ such that $w^{\alpha}=w^{\prime\prime}$.
(2.3) For every $\delta\in\Lambda_{n}^{m}$ such that $v_{n}^{m}(\delta)=1$, $w^{\delta}(\alpha)=2$ for every $w^{\delta}$. Define $v_{n}^{m+1}(\alpha)=2$. Clearly, $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$.
(3) $\alpha=\beta\#\gamma$ such that $^{\prime S4}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{1\}$ (for some $\in\{\land$$\to\}$). There are two subcases to analyze:
(3.1) $\alpha=\beta\to\gamma$. Then $v_{n}^{m}(\beta)=2$ and $v_{n}^{m}(\gamma)=1$. Define $v_{n}^{m+1}(\alpha)=1$. Then $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$ such that $w^{\alpha}=w^{\gamma}$. Indeed, since $w^{\gamma}(\beta)=2$ and $w^{\gamma}(\gamma)=0$ then $w^{\gamma}(\alpha)=0$.
(3.2) $\alpha=\beta\land\gamma$. Then $v_{n}^{m}(\beta)=1$ and $v_{n}^{m}(\gamma)\in\mathcal{D}$ or vice versa. Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ (where $w=w^{\beta}$ or $w=w^{\gamma}$) such that $w(\beta)=0$ or $w(\gamma)=0$ and so $w(\alpha)=0$. Define $v_{n}^{m+1}(\alpha)=1$. Then $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$ such that $w^{\alpha}=w^{\beta}$ or $w^{\alpha}=w^{\gamma}$.
(4) $\alpha=\Box\beta$ such that $\Box^{S4}v(\beta)=\{2\}$ or $\alpha=\beta\#\gamma$ such that $^{\prime S4}(v_{n}^{m}(\beta),v_{n}^{m}(\gamma))=\{2\}$ (for some $#\in\{\vee,\land,\to\}$). Define $v_{n}^{m+1}(\alpha)=2$. It is clear that $v_{n}^{m+1}\in PLV^{\prime}(\Lambda_{n}^{m+1})$.
We have shown in cases (1)-(4) how to extend the domain of $v_{n}^{m}$ to the additional formula $\alpha_{\omega\cdot(n+1)+m}$, showing that the resulting function $v_{n}^{m+1}$ is in $PLV^{\prime}(\Lambda_{n}^{m+1})$, and $\tilde{v}_{p}\subseteq v_{i}^{j}\subseteq v_{k}^{r}$, for every $(i,j)\leq(k,r)\leq(n,m+1)$. That is, we show how to increase the superscript $m$. In order to increase the subscript $n$, let $v_{n+1}=v_{n+1}^{0}:=\bigcup_{m\in\omega}v_{n}^{m}$. By the procedure described above, it is immediate to see that $v_{n+1}^{0}$ is in $PLV^{\prime}(\bigcup_{m\in\omega}{\Lambda}_{n}^{m})=PLV^{\prime}({\Lambda}_{n+1})=PLV^{\prime}({\Lambda}_{n+1}^{0})$, and $\tilde{v}_{p}\subseteq v_{i}^{j}\subseteq v_{k}^{r}$, for every $(i,j)\leq(k,r)\leq(n+1,0)$. By repeating the process, it is possible to define $v_{n}^{m}$ for every $(n,m)\in\omega\times\omega$ with the required properties. Finally, let $v:=\bigcup_{n\in\omega}v_{n}$. From the manner in which the construction was made, it follows that $v$ is in $PLV^{\prime}(\bigcup_{n\in\omega}{\Lambda}_{n})=PLV^{\prime}(For(\Sigma))$, and $\tilde{v}_{p}\subseteq v$. By the first part of the proof, we conclude that $v\in\mathcal{L}^{\prime}_{S4}$. ∎
@@
Let $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. Then, $\tilde{v}_{p}\in PV(\Lambda)$. Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$. Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},w,\alpha)$. Let $\tilde{w}_{p}=w_{|\Lambda}$. By Lemma 6.13, $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$. Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$. By Lemma 6.18, there exists $v^{\prime}\in\mathcal{L}^{\prime}_{S4}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$. Hence, $\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},v^{\prime},\alpha)$. This shows that
$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}^{4}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$}\big)\}.\end{array}$ From this, it follows that $PLV(\Lambda)=PLV^{\prime}(\Lambda)$, by Proposition 6.10. ∎
@@
7 First steps towards an RNmatrix for $\mathbf{IPL}$
As mentioned in Section 1, Kearns’s level valuations fail to provide an effective decision procedure for S4, while Grätz method produces an algorithm to decide that logic. Clearly, the composition of Gödel’s translation from $\mathbf{IPL}$ to S4 with Grätz’s decision procedure for S4 naturally produces a decision procedure for $\mathbf{IPL}$.
Let $\Sigma=(\{\neg,\Box\},\{\land,\vee,\to\})$ be the full signature of S4, and let $\Omega=(\{\neg\},\{\land,\lor,\rightarrow\})$ be the signature for intuitionistic propositional logic $\mathbf{IPL}$.
@@ 7 First steps towards an RNmatrix for $\mathbf{IPL}$
The complexity $\textsf{c}(\alpha)$ of a formula $\alpha\in For(\Omega)$ is defined as follows: $\textsf{c}(p)=0$ if $p$ is a propositional variable; $\textsf{c}(\neg\alpha)=\textsf{c}(\alpha)+1$; and $\textsf{c}(\alpha\#\beta)=\textsf{c}(\alpha)+\textsf{c}(\beta)+1$, for $#\in\{\to,\vee,\land\}$.
Recall the Gödel-Tarski-McKinsey Box translation $t:For(\Omega)\rightarrow For(\Sigma)$ defined as follows:
  1. i)
    $t(p)=\Box p$, if $p$ is a propositional variable;
  2. ii)
    $t(\neg\alpha)=\Box\neg t(\alpha)$;
  3. iii)
    $t(\alpha\#\beta)=\Box(t(\alpha)\#t(\beta))$, where $#\in\{\rightarrow,\lor,\land\}$.
It is widely known that this translation is faithful (see [25]; for a more recent proof see [26]), that is:
Theorem 7.2 ([25]).
Let $\varphi$ be a formula in $For(\Omega)$. Then: $\varphi$ is valid in $\mathbf{IPL}$ if and only if $t(\varphi)$ is valid in S4.
As seen in Section 6, Grätz’s results and algorithm can be extended to S4 defined over $\Sigma$, now based on $\mathcal{M}^{\prime}_{S4}$. Then, by Theorem 7.2 and by Theorem 6.20, a formula $\alpha$ in $For(\Omega)$ is valid in $\mathbf{IPL}$ iff the reduced truth-table obtained by taking the partial valuations in $\mathcal{M}^{\prime}_{S4}$ for the set of subformulas of the formula $t(\alpha)$ in $For(\Sigma)$ shows that $t(\alpha)$ is a valid formula in S4. This produces obviously a decision procedure for $\mathbf{IPL}$, based on an RNmatrix for S4 and a translation between $\mathbf{IPL}$ and S4. It is a challenging question how to define a direct decision procedure for $\mathbf{IPL}$ based on a finite-valued RNmatrix for $\mathbf{IPL}$, with level valuations (and with an associate truth-tables decision procedure) specifically designed for this logic. Moreover, the soundness and completeness of the RNmatrix (and the decision procedure obtained from it) should not rely on the corresponding soundness and completeness results of Gödel’s [11] and Grätz’s constructions. A solution to this question will be given in the next two sections.
8 An RNmatrix for $\mathbf{IPL}$
This section describes a new semantics for $\mathbf{IPL}$ based on a 3-valued RNmatrix with a very intuitive interpretation. We obtain this semantics by abstracting the composed procedure for $\mathbf{IPL}$ mentioned at the end of the previous section but defined in a direct way. That is, we describe a ‘pure’ RNmatrix for $\mathbf{IPL}$ (i.e., one that does not depend on any other construction), which turns out to be a suitable semantics for $\mathbf{IPL}$. Moreover, Section 9 presents an algorithm for deciding validity in $\mathbf{IPL}$ by removing spurious rows from the truth-tables generated by the procedure. Hence, the soundness and completeness of the semantics and the associated decision procedure are proven independently of Gödel’s 1933 results (proved in [25]) and those in [12].
8.1 Inducing the Nmatrix $\mathcal{M}_{IPL}$ for $\mathbf{IPL}$
The truth-tables for intuitionistic logic can be defined based on the Box translation $t$ described above, combined with a semi-translation $t_{0}$ to be defined now.
@@ 8.1 Inducing the Nmatrix $\mathcal{M}_{IPL}$ for $\mathbf{IPL}$
The semi-translation function $t_{0}:For(\Omega)\rightarrow For(\Sigma)$ is defined as follows:
@@
    $t_{0}(\alpha\#\beta)=t(\alpha)\#t(\beta)$, where $#\in\{\rightarrow,\lor,\land\}$.
Lemma 8.2.
For every $\alpha\in For(\Omega)$, $t(\alpha)=\Box(t_{0}(\alpha))$.
Proof.
Straightforward, by induction on the complexity of $\alpha$. ∎
@@
Observe that the interpretation of $\Box$ in $\mathcal{M}^{\prime}_{S4}$ is deterministic (see Table 2). Let us denote by $\Box^{4}(a)$ the only element of ${\Box^{\prime}}^{S4}(a)$, for $a\in\mathcal{V}$. Hence, if $v$ is a valuation in $\mathcal{M}^{\prime}_{S4}$ then $v(\Box\alpha)=\Box^{4}(v(\alpha))$, for every $\alpha\in For(\Sigma)$. By Lemma 8.2, $v(t(\alpha))=\Box^{4}(v(t_{0}(\alpha)))$, for every $\alpha\in For(\Omega)$.
To check the validity in $\mathbf{IPL}$ of $\alpha\in For(\Omega)$, one must construct a truth-table in $\mathcal{M}^{\prime}_{S4}$ for $t(\alpha)\in For(\Sigma)$. Assume that the set of subformulas of $\alpha$ is arranged as a sequence $\alpha_{1},\ldots,\alpha_{n}=\alpha$ in non-decreasing order of complexity (the complexity of formulas in $For(\Omega)$ is defined as usual, see Definition 7.1 above).
But then, the columns of the truth-table for $t(\alpha)$ in $\mathcal{M}^{\prime}_{S4}$ can be organized in pairs of formulas of the form $\langle t_{0}(\alpha_{1});t(\alpha_{1})\rangle,\ldots,\langle t_{0}(\alpha_{n});t(\alpha_{n})\rangle=\langle t_{0}(\alpha);t(\alpha)\rangle$. Indeed, the sequence of formulas $t_{0}(\alpha_{1}),t(\alpha_{1}),t_{0}(\alpha_{2}),\ldots,t_{0}(\alpha_{n}),t(\alpha_{n})=t(\alpha)$ in $For(\Sigma)$ induced by the sequence of pairs of formulas is the sequence of formulas that can be considered to construct a truth-table for $t(\alpha)$. [^5] Rigorously speaking, if $\alpha$ depends on the propositional variables $p_{1},\ldots,p_{k}$ then usually it is considered $\alpha_{i}=p_{i}$ (for $1\leq i\leq k$), and then $t(\alpha_{i})=\Box p_{i}$ is placed at a latter stage, say $\alpha_{i+k}$. Hence, the first $2k$ columns of a ‘standard’ table for $t(\alpha)$ starts with formulas $p_{1},\ldots,p_{k},\Box p_{1},\ldots\Box p_{k}$ instead of $p_{1},\Box p_{1},p_{2},\Box p_{2},\ldots p_{k},\Box p_{k}$. After this, the pairs produce the final columns of a table for $t(\alpha)$. For instance, in order to check $\alpha=p\vee\neg p$ it is considered the sequence of pairs $\langle p;\Box p\rangle$, $\langle\neg\Box p;\Box\neg\Box p\rangle$, $\langle\Box p\lor\Box\neg\Box p;\Box(\Box p\lor\Box\neg\Box p)\rangle$ based on the sequence $p,\neg p,p\vee\neg p$ for $\alpha$. Observe that $t(\alpha)=\Box(\Box p\lor\Box\neg\Box p)$.
This suggests considering for $\mathbf{IPL}$ an Nmatrix (derived from $\mathcal{M}^{\prime}_{S4}$) with domain $\left\{\langle c;\Box^{4}(c)\rangle\mid c\in\mathcal{V}\right\}=\{\langle 2;2\rangle,\langle 1;0\rangle,\langle 0;0\rangle\}$ and multioperations defined as follows:
  1. 1.
    $\neg^{IPL}(\langle a;b\rangle)=\left\{\langle c;\Box^{4}(c)\rangle\mid c\in{\neg^{\prime}}^{S4}(b)\right\}$
  2. 2.
    $^{IPL}(\langle a;b\rangle,\langle c;d\rangle)=\left\{\langle e;\Box^{4}(e)\rangle\mid e\in{\#^{\prime}}^{S4}(b,d)\right\}$
where $#\in\{\rightarrow,\lor,\land\}$.
Let us now consider a label for the truth-values above as follows: $\mathbf{T}:=\langle 2;2\rangle$, $\mathbf{U}:=\langle 1;0\rangle$ and $\mathbf{F}:=\langle 0;0\rangle$. The observations above lead us to the following:
Definition 8.4 (Intuitionistic Nmatrix $\mathcal{M}_{IPL}$).
The intuitionistic Nmatrix over signature $\Omega$ is given by $\mathcal{M}_{IPL}=\langle V_{IPL},\mathcal{D}_{IPL},\mathcal{O}_{IPL}\rangle$, where $V_{IPL}=\{\mathbf{T},\mathbf{U},\mathbf{F}\}$, $\mathcal{D}_{IPL}=\{\mathbf{T}\}$, and $\mathcal{O}_{IPL}(\#)=\#^{IPL}$ for each connective #\# in $\Omega$ is defined as in Figure 3.
Figure 3: Tables for $\mathcal{M}_{IPL}$.
Observe that, in a given valuation over $\mathcal{M}_{IPL}$, a formula is false when receives a non-designated value, $\mathbf{U}$ or $\mathbf{F}$. By looking at the tables of the Nmatrix, apparently both values have the same behavior. But it is not the case.
The first thing to note in Figure 3 is that the truth-value $\mathbf{U}$, which intuitively means that the epistemic status of $\alpha$ is unknown, only appears in the definitions of $\neg^{IPL}$ and $\to^{IPL}$. As it is going to be more detailed in the completeness proof, the intended meaning is the following (see also Remark 8.15 below):
$v(\neg\alpha)=\mathbf{U}$
only happens when both $\neg\alpha$ and $\alpha$ are false; that is, when there is no proof of (or knowledge or information about) neither $\alpha$ nor $\neg\alpha$, hence excluded middle fails for $\alpha$ (intuitionistic, non-classical scenario);
$v(\neg\alpha)=\mathbf{F}$
means that $\neg\alpha$ is false, but $\alpha$ is true; that is, although $\neg\alpha$ is unknown, $\alpha$ is known, therefore excluded middle holds for $\alpha$ (classical scenario);
$v(\alpha\to\beta)=\mathbf{U}$
only happens when both $\alpha$ and $\alpha\to\beta$ are false; that is, there is no proof of (or knowledge or information about) neither $\alpha$ nor $\alpha\to\beta$, hence Peirce’s law $\alpha\vee(\alpha\to\beta)$ fails for $\alpha$ and $\alpha\to\beta$ (intuitionistic, non-classical scenario);
$v(\alpha\to\beta)=\mathbf{F}$
means that $\alpha\to\beta$ is false, but $\alpha$ is true; that is, although $\alpha\to\beta$ is unknown, $\alpha$ is known, therefore Peirce’s law holds for $\alpha$ and $\alpha\to\beta$ (classical scenario).
In turn, the behavior of conjunction and disjunction is ‘classical’ (in terms of designated/non-designated values), and that is to be expected: the inference rules for conjunction and disjunction are identical in classical and intuitionistic logic. Because of this, $\mathbf{U}$ never appears as a value assigned to a conjunction or a disjunction.
The third point to note is that, if we disregard the truth-value $\mathbf{U}$, the remaining Nmatrix corresponds exactly to classical logic. In fact, it is straightforward to see how classical logic can be fully recovered from the intuitionistic Nmatrix. This recovery can be achieved by refining $\mathcal{M}_{IPL}$ (in the sense of [3] and [4]) to the truth-values $\mathbf{T},\mathbf{F}$}. The resulting is an Nmatrix with zero degree of non-determinism, equivalent to the standard (deterministic) matrix for classical logic.
8.2 Level valuations for $\mathcal{M}_{IPL}$
The following step is the definition of a suitable set of valuations (level valuations) over the Nmatrix for $\mathbf{IPL}$ just defined. The construction of the sets of level valuations given by Kearns and Grätz is based on the idea of super designated truth-values, that is: designated values $a$ such that any element of $\tilde{\Box}(a)$ is still designated, where $\tilde{\Box}$ is the multioperator associated to $\Box$. Thus, if a formula gets a designated value for every $n$th-level valuation, it must receive a super-designated value in every $(n+1)$th-level valuation. This guarantees the validity of the Necessitation inference rule, level by level. In this way, if a formula is provable in S4 with $n$ applications of the necessitation rule, then it is valid in $\mathcal{L}^{S4}_{n}$. From this, soundness of syntactical proofs in S4 w.r.t. the RNmatrix semantics follows easily. Necessitation is usually called a global inference rule: in order to be sound, the hypothesis of the rule must be provable, not merely assumed. However, $\mathbf{IPL}$ has no global inference rules in its usual Hilbert-style presentation (being Modus Ponens the only inference rule, which is local in opposition to global).
Taking into account that we are abstracting the composition of the Box translation with Grätz’s truth-tables algorithm for S4, the set of $n$th-level valuations will be constructed based on partial level valuations (Definition 5.5) instead of the Kearns-like Definition 5.1 of level valuations. Thus, the requirement that every valuation (row) $v_{p}$ such $v_{p}(\alpha)=1$ must be supported by another valuation (row) $w_{p}$ with $w_{p}(\alpha)=0$ while preserving the values $2$ can be easily adapted to the present framework, by changing $1$, 0 and $2$ by $\mathbf{U}$, $\mathbf{F}$ and $\mathbf{T}$, respectively (see Definition 8.6 below).
However, as observed above, there is no counterpart to Necessitation rule in the usual Hilbert-style calculus for $\mathbf{IPL}$, therefore the level valuations seem, in principle, unrelated to syntactical proofs in $\mathbf{IPL}$. The key is to consider a standard sequent calculus presentation for $\mathbf{IPL}$, and so it can be established, in a similar way to the case of S4, a relationship between syntactical proofs in $\mathbf{IPL}$ and nthnth-level valuations: if a sequent is proved in $\mathbf{IPL}$ with a derivation of height $n$ then it is valid w.r.t. nthnth-level valuations for $\mathbf{IPL}$ (see Proposition 8.11). Using this, the soundness of the sequent calculus for $\mathbf{IPL}$ w.r.t. its RNmatrix semantics can be easily obtained.
In addition, intuitionistic level valuations have a very interesting interpretation. Indeed, a suitable and very natural $3$-valued characteristic function associated to $\varphi$-saturated sets in $\mathbf{IPL}$ will be introduced in Definition 8.14. Hence, in Proposition 8.17 it will be proven that these characteristic functions correspond to level valuations.
The relationship between level valuations and $3$-valued characteristic functions of $\varphi$-saturated sets also has an interesting consequence in terms of complexity: at level 0 (and hence at any level) the valuations can only assign the classical truth-values $\mathbf{T}$ and $\mathbf{F}$ to propositional variables. Indeed, atomic formulas contain neither negation nor implication connectives, so it makes no sense to assign them the value $\mathbf{U}$, which, as discussed above, is exclusively related to the failure of the law of excluded middle and Peirce’s law. This feature will reduce the complexity of the algorithm, as we will be discuss in Section 10.